# Are foundation models for computer vision good conformal predictors?

### v1

@@ 4.2.2 Impact under Distribution Shifts
The theoretical guarantees of the coverage for conformal prediction hold under the hypothesis that the calibration set and the test set are drawn from the same distribution, i.e., data exchangeability assumption. Thus, in this section, we analyze the impact of having calibration sets that present distributional shifts with regard to the testing set.
@@
APS [40] exhibits strong robustness against large distributional shifts, at the cost of substantially degrading efficiency. One would expect that adaptive CP methods, such as APS and RAPS, somehow mitigate domain shifts due to their adaptive nature. Nevertheless, Fig. 4 reveals several interesting observations, which contradict this intuition. First, we can observe that, when resorting to APS as CP method the coverage gap is consistently satisfied (or nearly satisfied) across all domain shifts and models (Fig. 4, middle). In contrast, RAPS generally shows very similar performances compared to the straighforward LAC, obtaining lower marginal coverage under several models and domains, and substantially lower than APS. To understand this phenomenon, we now study how the set sizes evolve across domains for the different methods (Fig. 4, left). We can easily observe that (i) APS yield the largest conformal sets across ImageNet domains, regardless of the model, and (ii) APS experiences the largest set increases when the complexity of the domain grows. Thus, as exposed in the previous Section, APS satisfied the marginal coverage by substantially including more predicted classes, therefore increasing conformal set sizes.
@@
APS and RAPS are adaptive methods that produce similar minimum class-conditional coverage, as exhibited in ImageNet-A (Fig. 4, right). However, they showcase significant differences in the average coverage gap. In the following, we explore this phenomenon in detail. More concretely, we depict in Fig. 5 the distribution of the conditional class coverage values obtained by APS and RAPS on ImageNet-A. Interestingly, while both approaches see their minimum class-conditional coverage decrease, their distributions are completely different. Indeed, APS distributions exhibit a Gaussian shape, with a decreasing number of categories presenting lower conditional coverage as they separate from 1-$\alpha$. In contrast, the distribution in RAPS exposes a significantly worse scenario, where class-conditional coverages are almost uniformly spread, with a non-negligible amount of classes below the expected coverage, i.e., $1-\alpha$.
Another interesting, and valuable observation is that the low performance of all CP methods is magnified when coupled with VICReg (i.e., larger set sizes and lower coverage), suggesting that this family of foundation models may suffer more under distribution shifts.
@@ 4.2.3 Does calibration affect Conformal Prediction?
Confidence calibration is a popular strategy to improve the uncertainty estimates of deep models. These techniques, which can be either added as a post-processing step [36, 13, 24] or integrated as a training regularizer [16, 5, 32], adapt the model softmax predictions to reflect their actual performance accurately. The exponentially growing adoption of vision and vision-language foundation models in critical areas makes integrating confidence calibration a natural progression, as evidenced by recent works [30, 55, 50]. Thus, in this section, we investigate this important issue, as the relationship between calibration and conformal prediction in vision foundation models remains largely unexplored. Specifically, we examine whether calibrating these models with the popular Temperature Scaling (TS) [13] affects the conformal performance of fixed and adaptive CP methods. Particularly, we apply TS to the ImageNet results in the general case (Section 4.2.1). As a proper validation set is not available, we evaluate the CP performance over a set of $T$ values (14 values from $0.85$ to $2$), and found that $T=1.1$ typically yield well-calibrated models [^1] Note that our goal is not to obtain the best-calibrated model but evaluate the impact of calibration on the conformal predictions..
Calibration decreases the efficiency of conformal prediction sets but increases the minimum class-conditional coverage, improving the coverage gap. Confidence calibration typically smooths the distribution of the class softmax scores, which results in less confident predictions. Consequently, the dominant value in the predicted softmax vector is typically lower in calibrated models. Nevertheless, since CP methods provide theoretical guarantees (under the data exchangeability assumption) [52] to satisfy the target $1-\alpha$ of marginal coverage, these changes in the softmax distributions affect the conformalization obtained pre-TS. In particular, in Table 2, we present the results for the average set size and minimum class-conditional coverage before and after scaling the logits with TS ($T=1.1$). Furthermore, we include the ECE values to verify that model calibration has improved. These results show that if the model is calibrated, its prediction set size tends to be larger, particularly for adaptive CP approaches (RAPS, and more specifically, APS). Fig. 7 further delves into these results, where we plot the distribution of differences between the set size of samples before and after calibration (i.e., a point in the distribution is $\mathcal{C}(x_{i})-\mathcal{C}(x_{i}^{TS})$). We can identify that the overall larger size in APS is caused by a consistent efficiency degradation across samples and not a few atypical cases with large conformal sets.
@@
We observe that class-conditional coverage is typically improved on calibrated models, particularly when they are conformalized by adaptive CP methods. Despite being marginal, these improvements suggest that calibrating vision foundation models decrease efficiency while marginally enhancing class-conditional coverage.
The best coverage gap obtained by APS [40] approximately aligns with the coverage gap achieved at the optimal calibration point, whereas RAPS [1] coverage gap strongly differs. Fig. 6 shows the impact of the temperature $T$ on the different CP metrics on CLIP conformalized predictions. An interesting observation is that at the optimal temperature value $T=1.1$, the CovGap of APS is very close to its optimal point [^2] Results on Appendix D show similar behavior for other models. (0.0567 vs 0.0561). In contrast, while the RAPS coverage gap at optimal calibration is 0.0701, it decreases to 0.0642 as $T$ increases. This suggests that we could anticipate the best coverage gap for APS by assessing the model calibration performance, which will likely be near optimal calibration. On the downside, APS conformal sets efficiency is degraded. As $T$ increases, the softmax distributions get smoother, explaining the degradation in efficiency, which is even more drastic for APS, whose set size monotonically increases with $T$. We concede that, whereas APS minimizes the coverage gap, it does so at the expense of increasing the conformal set size, particularly compared to other methods. However, we believe that increasing the size, up to some extent, while maintaining a better coverage gap is preferable in critical decision systems.
@@ 4.2.4 Effect on few-shot adapted models
Adapting zero-shot CLIP models for downstream tasks in a few-shot labeled regime is becoming increasingly popular. These strategies can be mainly categorized into prompt learning [58, 14, 57], which optimize the set of text prompts given to the text encoder, and Adapters [56, 25, 19, 45, 58, 14], where only a limited set of learnable parameters atop embedding representations is updated. Thus, questioning whether adapting these models hinges the performance of CP methods is of paramount importance, as it addresses a foundational aspect of effectively integrating uncertainty quantification with modern architectures. To assess the impact of CLIP adaptation in conformal prediction, we resort to representative methods of the few-shot adaptation families presented above, and follow standard adaptation and evaluation protocols in the literature [45]: adapt CLIP on few-shot samples ($M=16$) from ImageNet, and evaluate on ImageNet and its variants.
@@
ID ZS 2.81 10.07 3.23 0.086 0.069 0.084
ZSLP 2.17 6.60 2.45 0.076 0.060 0.074
CLAP [45] 2.15 6.83 2.45 0.077 0.060 0.074
CoOp [58] 2.44 6.32 2.79 0.080 0.066 0.080
KgCoOp [14] 2.47 7.18 2.84 0.081 0.065 0.081
@@
Table 3: Set size and CovGap for CLIP (ViT-B backbone) on in-distribution (ImageNet) and out-of-distribution (averaged over ImageNet versions) across 5 few-shot adaptation approaches.
Adapters are generally better candidates for Conformal Prediction than Prompt Learning methods. Table 3 reports the results of adapting CLIP with different approaches. First, we observe that, across both ID and OOD scenarios, Adapters (i.e., ZSLP and CLAP) typically yield smaller, or comparable, conformal set sizes than prompt learning (i.e., CoOp and KgCoOp), while always providing a lower CovGap, even compared to the ZS predictions. Furthermore, and more interestingly, conformalizing the predictions of all adaptation methods with APS in the OOD scenario improves the set efficiency compared to the ZS baseline. In contrast, LAC and RAPS typically increase the set size. These observations are related to the findings from the previous section, which suggested that better-calibrated models lead to larger conformal sets. Indeed, recent evidence [30] demonstrated that few-shot CLIP adaptation methods deteriorate the confidence estimates compared to ZS predictions, which, following our observations, should result in smaller conformal sets, as validated in Table 3.
@@ 5 Conclusion
In this study, we aim at answering the question: Which CP method and model should I use, and what can I expect, in the era of vision foundation models? Our findings revealed that vision foundation models yield better conformal scores than their traditional pre-trained counterparts, with models integrating visual transformers outperforming those that use convolutional neural networks, particularly under domain shifts. Furthermore, we observed that several common situations encountered in practice (i.e., presence of distributional drifts and models undergoing confidence calibration) are indeed detrimental for some CP approaches. Interestingly, an adaptive CP method, i.e., APS, exhibited stronger robustness to these scenarios, particularly in terms of conditional coverage, but at the expense of degrading set efficiency. Last, under the popular few-shot CLIP adaptation setting, adapters have demonstrated to be a more promising alternative than Prompt Learning.
The final decision ultimately hinges on the requirements of each task. In certain fields, e.g., medical diagnosis, it may be preferable to maximize the conditional coverage across classes, even if it increases the conformal set size, as this minimizes the risk of critical errors. In contrast, in domains where errors are less consequential, maintaining smaller set sizes might be preferable to streamline decision-making and reduce computational demands. Thus, the choice depends on balancing the trade-offs between accuracy, usability, and the specific demands of the application.
@@ A Metrics
Set $\text{Set size}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{x\in\mathcal{D}_{\text{test}}}|\hat{C}(x)|$ (7)
$\text{Cov}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{x,y\in\mathcal{D}_{\text{test}}}\mathbbm{1}_{y\in\hat{C}(x)}$ (8)
@@ D Model performance for the optimal temperature
Figures 8 and 9 showcase the impact of temperature $T$ on the ECE, set size, minimum class-conditional coverage, and coverage gap across all foundation models. The observations on these figures strongly align with the findings on the main paper as, regardless of the model, APS yields optimal coverage gap close to the optimal temperature point. Indeed, looking at the different models (i.e., DINO-based, VICReg and CLIP), we can observe that the behaviour of RAPS (in terms of coverage gap), strongly varies across models. Additionally, Fig. 10 shows the evolution of $q_{\alpha}$, the threshold on the conformal scores $s$ before and after calibration. When applying temperature scaling, the distribution of softmax across classes approaches a uniform distribution, lowering the scores for the most likely classes. Intuitively, this means that more classes will need to be included in the set to ensure coverage, leading to a decrease in threhsold $q_{\alpha}$. This decrease is consistent with the observed increase in set size, which can be seen in Fig. 11, showing the difference in set size when applying temperature scaling for APS on CLIP. Note that this trend holds across models, as a similar behaviour for DINO is observed in Fig. 7 in the main paper. Note that both Figures 7 and 11 only show cases where the set sizes are different between the calibrated and uncalibrated model.
@@
E Domain Shift
Table 8 reports the numerical values for Fig. 4, showcasing APS’s strong performance in terms of marginal and conditional coverage under distribution shift, at the cost of decreasing the set efficiency. Fig. 12, Fig. 13, and Fig. 14 depict the distribution of class-conditional coverage for DINOv2-B, VICReg (ResNet 50x2), and CLIP, respectively. The kernel density estimate plots show the class-conditional coverage for a distribution shift with ImageNet-A (top), ImageNet-R (middle), and ImageNet-V2 (bottom), for APS (left) and RAPS (right). These results clearly demonstrate the stronger resistance to distribution shift for APS compared to RAPS, even when both methods show a minimum class-conditional coverage of 0.
F Few-Shot Adaptation
We present in Table 9 a detailed version of Table 3, where the OOD section (in Table 3) corresponds to the average across all ImageNet variants.

### v2

@@ 4.2.2 Impact under Distribution Shifts
The theoretical guarantees of the coverage for conformal prediction hold under the hypothesis that the calibration set and the test set are drawn from the same distribution, i.e., data exchangeability assumption. In this section, we analyze the impact of having calibration sets that present distributional shifts with regard to the testing set.
@@
APS [48] exhibits strong robustness against large distributional shifts, at the cost of substantially degrading efficiency. One would expect that adaptive CP methods, such as APS and RAPS, somehow mitigate domain shifts due to their adaptive nature. Nevertheless, Fig. 4 reveals several interesting observations, which contradict this intuition. First, we can observe that, when resorting to APS as CP method the coverage gap is consistently satisfied (or nearly satisfied) across all domain shifts and models (Fig. 4, middle). In contrast, RAPS generally shows very similar performances compared to LAC, obtaining lower marginal coverage under several models and domains, and substantially lower than APS. To understand this phenomenon, we now study how set sizes evolve across domains for the different methods (Fig. 4, left). We can easily observe that (i) APS yield the largest conformal sets across ImageNet domains, regardless of the model, and (ii) APS experiences the largest set increases when the complexity of the domain grows. Thus, as exposed in the previous Section, APS satisfied marginal coverage by substantially including more predicted classes, therefore increasing conformal set sizes.
@@
APS and RAPS are adaptive methods that produce similar minimum class-conditional coverage, as exhibited in ImageNet-A (Fig. 4, right). However, they showcase significant differences in the average coverage gap. In the following, we explore this phenomenon in detail. More concretely, we depict in Fig. 5 the distribution of the conditional class coverage values obtained by APS and RAPS on ImageNet-A for CLIP (more datasets and models in Appendix F). Interestingly, while both approaches see their minimum class-conditional coverage decrease, their distributions are completely different. Indeed, APS distributions exhibit a Gaussian shape, with a decreasing number of categories presenting lower conditional coverage as they separate from 1-$\alpha$. In contrast, the distribution in RAPS exposes a significantly worse scenario, where class-conditional coverages are almost uniformly spread, with a non-negligible amount of classes below the expected coverage, i.e., $1-\alpha$.
Another interesting and valuable observation is that the low performance of all CP methods is magnified when coupled with VICReg (i.e., larger set sizes and lower coverage), suggesting that this family of foundation models may suffer more under distribution shifts.
@@ 4.2.3 Does calibration affect Conformal Prediction?
Confidence calibration is a popular strategy to improve the uncertainty estimates of deep models. These techniques, which can be either added as a post-processing step [44, 16, 29] or integrated as a training regularizer [19, 5, 38], adapt the model softmax predictions to reflect their actual performance accurately. The exponentially growing adoption of vision and vision-language foundation models in critical areas makes integrating confidence calibration a natural progression, as evidenced by recent works [36, 66, 59]. Thus, in this section, we investigate this important issue, as the relationship between calibration and conformal prediction in vision foundation models remains largely unexplored. Specifically, we examine whether calibrating these models with the popular Temperature Scaling (TS) [16] affects the conformal performance of fixed and adaptive CP methods. Particularly, we apply TS to the ImageNet results in the general case (Section 4.2.1). As a proper validation set is not available, we evaluate the CP performance over a set of $T$ values (14 values from $0.85$ to $2$), and found that $T=1.1$ typically yielded well-calibrated models [^1] Note that our goal is not to obtain the best-calibrated model but evaluate the impact of calibration on CP. Furthermore, this value aligns with existing works [34, 24] using TS for similar datasets, e.g., TinyImageNet..
Our observations suggest that calibration decreases the efficiency of CP sets but increases the minimum class-conditional coverage, improving the coverage gap.
Confidence calibration typically smooths the distribution of the class softmax scores, which results in less confident predictions. Consequently, the dominant value in the predicted softmax vector is typically lower in calibrated models. Nevertheless, since CP methods provide theoretical guarantees (under the data exchangeability assumption [61]) to satisfy the target marginal coverage of $1-\alpha$, these changes in the softmax distributions affect the conformalization obtained pre-TS. We present in Table 2 the results for the average set size and minimum class-conditional coverage before and after scaling the logits with TS. Furthermore, we include the Expected Calibration Error (ECE) values to verify that model calibration has improved. These results show that if the model is calibrated, its prediction set size tends to be larger, particularly for adaptive CP approaches (RAPS, and more specifically, APS). Fig. 7 further delves into these results, where we plot the distribution of differences between the set size of samples before and after calibration (i.e., a point in the distribution is $\mathcal{C}(x_{i})-\mathcal{C}(x_{i}^{TS})$). We can identify that the overall larger size in APS is caused by a consistent efficiency degradation across samples and not a few atypical cases with large conformal sets. We also observe that class-conditional coverage is typically improved on calibrated models, particularly when they are conformalized by adaptive CP methods. Thus, these results suggest that calibrating vision foundation models decreases efficiency while marginally enhancing class-conditional coverage, particularly on adaptive CP.
@@
The best coverage gap obtained by APS [48] approximately aligns with the coverage gap achieved at the optimal calibration point, whereas RAPS [1] coverage gap strongly differs. Fig. 6 shows the impact of the temperature $T$ on the different CP metrics on CLIP conformalized predictions. An interesting observation is that at the optimal temperature value $T=1.1$, the CovGap of APS is very close to its optimal point [^2] Results on Appendix D show similar behavior for other models. (0.0567 vs 0.0561). In contrast, while the RAPS coverage gap at optimal calibration is 0.0701, it decreases to 0.0642 as $T$ increases. This suggests that we could anticipate the best coverage gap for APS by assessing the model calibration performance, which will likely be near optimal calibration. On the downside, APS conformal sets efficiency is degraded. As $T$ increases, the softmax distributions get smoother, explaining the degradation in efficiency, which is even more drastic for APS, whose set size monotonically increases with $T$. We concede that, whereas APS minimizes the coverage gap, it does so at the expense of increasing the conformal set size. However, we believe that increasing the size, up to some extent, while improving coverage gap is preferable in critical decision systems.
@@ 4.2.4 Effect on few-shot adapted models
Adapting zero-shot CLIP models for downstream tasks in a few-shot labeled regime is becoming increasingly popular in VLMs. These strategies can be mainly categorized into Prompt Learning [70, 17, 69], which optimize the set of text prompts given to the text encoder, and Adapters [67, 30, 23, 53, 70, 17], where only a limited set of learnable parameters atop embedding representations is updated. Thus, questioning whether adapting these models hinges the performance of CP methods is of paramount importance, as it addresses a foundational aspect of effectively integrating uncertainty quantification with modern architectures. To assess the impact of CLIP adaptation in conformal prediction, we resort to representative methods of the few-shot adaptation families presented above, and follow standard adaptation and evaluation protocols in the literature [53]: (i) for out-of distribution (OOD), we adapt CLIP on few-shot samples ($M=16$) from ImageNet, and evaluate on ImageNet and its variants, and (ii) for in distrbution (ID) we adapt on few-shot samples and evaluate on the validation set from the same dataset.
@@
ID ZS 5.29 6.98 6.49 0.114 0.094 0.102
ZSLP 2.09 3.43 2.53 0.087 0.060 0.064
CLAP [53] 2.13 3.52 2.57 0.088 0.060 0.065
CoOp [70] 2.07 2.87 2.47 0.083 0.059 0.067
KgCoOp [17] 2.10 3.11 2.53 0.085 0.059 0.065
@@
Table 3: Few-shot adaptation (16-shots). Set size and CovGap for CLIP (ViT-B backbone) on in-distribution (average over 11 datasets) and out-of-distribution (average over ImageNet versions). Results with 4-shots are provided in Appendix, Table 10.
Few-shot VLM adaptation render pre-trained models better conformalized in the ID scenario (Table 3). First, we observe that, across both ID and OOD scenarios, both Adapters and Prompt Learning yield smaller conformal set sizes and higher Coverage gap than zero-shot CLIP, regardless of the CP method, with Prompt Learning yielding slightly better performances. In contrast, on the OOD scenario, only APS consistently enhances the set efficiency of the ZS model, with all the few-shot adaptation methods obtaining scarce coverage gap improvements across CP methods. These observations are related to the findings from the previous section, which suggested that better-calibrated models lead to larger conformal sets. Indeed, recent evidence [36] demonstrated that few-shot CLIP adaptation methods deteriorate the confidence estimates compared to ZS predictions, which, following our observations, should result in smaller conformal sets, as validated in Table 3. Further details are provided in Appendix G.
@@ 5 Conclusion
In this study, we aimed at answering the question: Which CP method and model should I use, and what can I expect, in the era of vision foundation models? Our findings revealed that vision foundation models yield better conformal metrics than their traditional pre-trained counterparts, with models integrating visual transformers outperforming those that use convolutional neural networks, particularly under domain shifts. Furthermore, we observed that several common situations encountered in practice (i.e., presence of distributional drifts and models undergoing confidence calibration) are indeed detrimental for some CP approaches. Interestingly, an adaptive CP method, i.e., APS, exhibited stronger robustness to these scenarios, particularly in terms of conditional coverage, but at the expense of degrading the set efficiency.
@@ A Metrics
Set $\text{Set size}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{\mathbf{x}\in\mathcal{D}_{\text{test}}}|\hat{C}(\mathbf{x})|$ (7)
$\text{Cov}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{\mathbf{x},y\in\mathcal{D}_{\text{test}}}\mathbbm{1}_{y\in\hat{C}(\mathbf{x})}$ (8)
@@ D Model performance for the optimal temperature
Fig. 8, 9, 10, 11, 12, and 13 showcase the impact of temperature $T$ on the ECE, set size, minimum class-conditional coverage, and coverage gap across all foundation models. The observations on these figures strongly align with the findings on the main paper as, regardless of the model, APS yields optimal coverage gap close to the optimal temperature point. Indeed, looking at the different models (i.e., DINO-based, VICReg and CLIP), we can observe that the behaviour of RAPS (in terms of coverage gap), strongly varies across models. Additionally, Fig. 14 shows the evolution of $q_{\alpha}$, the threshold on the conformal scores $s$ before and after calibration. When applying temperature scaling, the distribution of softmax across classes approaches a uniform distribution, lowering the scores for the most likely classes. Intuitively, this means that more classes will need to be included in the set to ensure coverage, leading to a decrease in threhsold $q_{\alpha}$. This decrease is consistent with the observed increase in set size, which can be seen in Fig. 15, showing the difference in set size when applying temperature scaling for APS on CLIP. Note that this trend holds across models, as a similar behaviour for DINO is observed in Fig. 7 in the main paper. Note that both Fig. 7 and 15 only show cases where the set sizes are different between the calibrated and uncalibrated model.
@@
E Conformal set size analysis
We show the difference between the conformal set size for APS when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{DINO-S}}$ in Fig. 20 and for $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{MetaCLIP}}$ in Fig. 19, similarly to what is shown in Fig. 4.
F Domain Shift
Table 8 reports the numerical values for Fig. 4, showcasing APS’s strong performance in terms of marginal and conditional coverage under distribution shift, at the cost of decreasing the set efficiency. Fig. 16, 17, and 18 depict the distribution of class-conditional coverage for DINOv2-B, VICReg (ResNet 50x2), and CLIP, respectively. The kernel density estimate plots show the class-conditional coverage for a distribution shift with ImageNet-A (top), ImageNet-R (middle), and ImageNet-V2 (bottom), for APS (left) and RAPS (right). These results clearly demonstrate the stronger resistance to distribution shift for APS compared to RAPS, even when both methods show a minimum class-conditional coverage of 0.
G Few-Shot Adaptation
We present in Table 9 and Table 11 a detailed version of Table 3 for 16 shots, where the OOD section (in Table 3) corresponds to the average across all ImageNet variants of Table 9, and the ID section corresponds to the average across all 11 datasets of Table 11. Table 10 shows in distribution results for adapting only 4 shots, where we observe slightly higher set size and coverage gap compare to the results for 16 shots.
@@ G Few-Shot Adaptation
(a) DINO-S
(b) DINO-B
(c) DINOv2-S
(d) DINOv2-B
(e) DINOv2-L
(f) DINOv2-G
Figure 10: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for DINO-S, DINO-B, DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G on CIFAR-10. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet 50)
(b) VICReg (ResNet 50x2)
(c) VICReg (ResNet 200x2)
(d) CLIP
(e) MetaCLIP
Figure 11: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for VICReg (ResNet 50), VICReg (ResNet 50x2), VICReg (ResNet 200x2), and CLIP on CIFAR-10. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) DINO-S
(b) DINO-B
(c) DINOv2-S
(d) DINOv2-B
(e) DINOv2-L
(f) DINOv2-G
Figure 12: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for DINO-S, DINO-B, DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G on CIFAR-100. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet 50)
(b) VICReg (ResNet 50x2)
(c) VICReg (ResNet 200x2)
(d) CLIP
(e) MetaCLIP
Figure 13: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for VICReg (ResNet 50), VICReg (ResNet 50x2), VICReg (ResNet 200x2), and CLIP on CIFAR-100. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
@@
Figure 19: $\text{ViT}_{\text{ImageNet}}$ vs $\text{ViT}_{\text{MetaCLIP}}$. Analyzing the difference in set size for APS between a $\text{ViT}_{\text{MetaCLIP}}$ and $\text{ViT}_{\text{ImageNet}}$. Equal set sizes not shown.
Figure 20: $\text{ViT}_{\text{ImageNet}}$ vs $\text{ViT}_{\text{DINO-S}}$. Analyzing the difference in set size for APS between a $\text{ViT}_{\text{DINO-S}}$ and $\text{ViT}_{\text{ImageNet}}$. Equal set sizes not shown.